# CMB delensing with deep learning

### v2

@@ 3.3 Loss function
3.4 Training and testing
Table 2: Adjustment and setting of hyperparameters in the U-Net++ architecture design. Prior values indicates that the optimum value is selected from the parameters of the preset value.
Hyperparameter Description Prior values Optimum
$\eta$ learning rate [$10^{-3},10^{-4},10^{-5}$] $10^{-4}$
$\omega$ weight decay [$10^{-4},10^{-5},10^{-6}$] $10^{-5}$
@@ 3.4 Training and testing
Figure 8: Loss function evolution for each network as a function of the number of epochs. From top to bottom, represent the results of training for T, Q, and U maps, respectively. The dark blue solid line indicates the training set loss function evolution and the light blue dashed line indicates the validation set loss function evolution.
@@
U-Net networks are predicated on fully convolutional networks (FCNs), which comprise a convolutional network and an inverse convolutional network. Thus, the heart of these networks is the convolutional layer, which involves convolving filters on the input data.
@@
In alignment with the works of Makinen et al. and Ni et al. (Makinen et al., 2021; Ni et al., 2022), we have fixed the number of convolutional kernels at the outset of the input to $32$. The kernel size determines the convolution’s field of view, which we have established at $3\times 3$. To achieve the requisite output dimensionality, we have employed “same" padding to manage sample boundaries in both convolutions and transpose convolutions. The stride parameter specifies the kernel’s traversal steps across the image. In our model, we have maintained the default settings of ${\rm stride}=1$ for convolutions and ${\rm stride}=2$ for transpose convolutions.
@@
The U-Net++ architecture was used to train the CMB delensing process via a set of lensed and unlensed CMB sky maps in an end-to-end fashion. Table 2 displays the hyperparameters’ specifics used in this network. The Adam optimizer was utilized in the analysis with the standard TensorFlow parameters (Reddi et al., 2019). The hyperparameters were meticulously fine-tuned for network optimization. The batch size and the initial number of convolutional filters are optimized to $64$, $128$ and $256$, respectively, being restricted by GPU memory. The number of epochs is fixed to $3000$; Table 2 illustrates the learning rate setup. Meanwhile, weight decay was examined with a list of previous values listed in Table 2. The optimized values of the initial number of convolution filters, learning rate, weight decay, and batch size were fixed at $32$, $10^{-4}$, $10^{-5}$, 16, respectively. The entire number of trainable parameters is $7.4\times 10^{7}$. The total number of trainable parameters is $9.04\times 10^{6}$. Since the CMB sky map data contains both positive and negative values due to temperature fluctuations, ReLU and LeakyReLU activation with alpha=1.0 was chosen to handle this relationship. Figure 8 illustrates the evolution of the loss function during the training of T, Q, and U maps as a function of the number of epochs. The figure shows the training results of T, Q, and U maps in a top-to-bottom order. The dark blue line indicates the evolution of the training set loss function, while the light blue dashed line represents the validation dataset’s loss function development.
4 Results and Analysis
In this section, we conduct a thorough analysis of the results, exploring them in depth from both the sky map and power spectrum perspectives.
4.1 Sky map analysis
Figure 9: The results of U-Net++ model and the residuals between predictions and true values are displayed. From left to right, the image columns correspond to the samples of sky patches related to T, Q and U fields. From top to bottom: the first row is the test image (lensed sky patches), the second row is the prediction of the test image (delensed sky patches), the third row is the true image (unlensed sky patches), and the last row is the error between the prediction and the true image (residuals between delensed and unlensed). A comparable map schematic can be observed in Figure 7, featuring an image size of $214.86~{}{\rm{deg}}^{2}$ and centered at $(l,b)=(78.75^{\circ},0^{\circ})$. The measurement units remains $mK$.
@@ 4.1 Sky map analysis
Based on the network parameter configuration described in the previous chapter, we have trained multiple datasets and performed prediction operations on the test set. Figure 9 shows the residual performance between the predicted and true values of U-Net++ model processing results. First, in the first row of the figure, we present a small-scale map sample from the test set as a basic reference. Second, in the second row, we show the corresponding small-scale map after delensing processing. Similarly, for comparison and verification, we also provide true small-scale unlensed map data. Finally, in the last row of the figure, to verify the effectiveness of delensing processing, we further show the specific small-scale map of the difference between the predicted and true values.
@@
SSIM is mainly used to analyze the structural information of two images. The closer the SSIM is to $1$, the more similar the structural information of the two images is. We have generally processed SSIM by setting the parameters $\alpha,\beta,\gamma$ to $1$. For the sky patches of two fields $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, we have carried out the following treatment:
$\begin{split}\text{SSIM}(\mathcal{F}_{1},\mathcal{F}_{2})&=[l(\mathcal{F}_{1},\mathcal{F}_{2})^{\alpha}\cdot c(\mathcal{F}_{1},\mathcal{F}_{2})^{\beta}\cdot s(\mathcal{F}_{1},\mathcal{F}_{2})^{\gamma}]\\ &=\frac{(2\bar{\mathcal{F}}_{1}\bar{\mathcal{F}}_{2}+c_{1})[2\text{CoV}({\mathcal{F}_{1},\mathcal{F}_{2}})+c_{2}]}{(\bar{\mathcal{F}}_{1}\bar{\mathcal{F}}_{2}^{2}+c_{1})[\sigma(\mathcal{F}_{1})^{2}+\sigma(\mathcal{F}_{1})^{2}+c_{2}]},\end{split}$ (17)
where $l(\mathcal{F}_{1},\mathcal{F}_{2}),c(\mathcal{F}_{1},\mathcal{F}_{2}),s(\mathcal{F}_{1},\mathcal{F}_{2})$ represent the brightness comparison, contrast comparison, and structure comparison between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, respectively. $\text{CoV}(\mathcal{F}_{1},\mathcal{F}_{2})$ represents the covariance between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, while $\sigma(\mathcal{F}_{1})$ and $\sigma(\mathcal{F}_{2})$ represent the standard deviation of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, respectively. $c1$ and $c_{2}$ are constants to prevent the numerator or denominator from being zero.
We have calculated the structural similarity between the test images (lensed sky patches) and the real images (unlensed sky patches), as well as between the predicted images (delensed sky patches) and the real images, which is shown in Figure 9. The specific results are listed in Table 3. We can see from the table that the SSIM values of delensed T, Q, U sky patches are almost close to 1, indicating that their structures are very similar to those of unlensed sky patches.
Table 3: Table of SSIM values between test images (lensed sky patches) and their predicted images (delensed sky patches) and real images (unlensed sky patches) for T, Q, and U.
[…]
(train, truth) 0.7542 0.7065 0.7141
(pred, truth) 0.9952 0.9933 0.9942
@@
$PSNR(\mathcal{F}_{1}-\mathcal{F}_{2})=10\cdot\log_{10}[\frac{\text{MAX}(\mathcal{F}_{1},\mathcal{F}_{2})}{\text{MSE}(\mathcal{F}_{1},\mathcal{F}_{2})}]\text{dB},$ (18)
@@
where $\text{MAX}(\mathcal{F}_{1},\mathcal{F}_{2})$ represents the maximum pixel difference between images $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, and $\text{MSE}(\mathcal{F}_{1},\mathcal{F}_{2})$ represents the mean squared error between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. Through this calculation, we can more comprehensively evaluate the degree of distortion at the pixel level in the image, and thus provide strong support for the comprehensive evaluation of image quality.
@@
PSNR stands as a simple yet widely utilized metric for assessing image quality, employing the decibel (dB) as its unit of measurement. Within this evaluative framework, a higher numerical value signifies a lower degree of image distortion, thereby indicating superior image quality. Akin to the computation of the SSIM metric, we have also conducted a corresponding analytical calculation of the PSNR, as shown in Table 4.
Table 4: Table of PSNR values between test images (lensed sky patches) and their predicted images (delensed sky patches) and real images (unlensed sky patches) for T, Q, and U.
[…]
(train, truth) 33.01 26.54 26.87
(pred, truth) 46.96 45.69 46.30
PSNR measures the degree of image distortion by calculating the pixel differences between images. According to the results presented in Table 4, we can clearly observe that the delensed sky patches, compared to the original lensed sky patches, exhibit a significant advantage in terms of the PSNR metric, with notably higher values. This indicates that the delensed sky patches are superior in preserving details and conveying information, effectively reducing the degree of image distortion.
Although SSIM and PSNR highly affirm the accuracy of the predicted data in numerical terms, they only quantitatively compare the pixel values of the test data with the truth data, without considering whether the predicted images have the same physical meaning or research value as the real images. Therefore, we need to conduct in-depth research on the angular power spectrum of delensed CMB maps.
4.2 Angular power spectrum analysis
In this subsection of the study, we conducted a comprehensive analysis of the angular power spectrum of the CMB temperature and polarization. Following multiple iterations of U-Net++ network training, we present the power spectra of CMB TT, BB, EE, and TE, as illustrated in Figures 10-13. The upper portion of the figures depicts the power spectra, with red indicating lensed, blue denoting unlensed, and orange representing delensed. The lower section illustrates the ratios between lensed, unlensed, and delensed components. Specifically, the pink color corresponds to $C_{\ell}^{delensed}/C_{\ell}^{lensed}$, the light blue color signifies $C_{\ell}^{delensed}/C_{\ell}^{unlensed}$, and the gray color denotes $Ratio=1$.

### v3

@@ III.3 Loss function
III.4 Training and testing
Table 1: Adjustment and setting of hyperparameters in the UNet++ architecture design. Prior values indicates that the optimum value is selected from the parameters of the preset value.
H-Param Description Prior values Optimum
$\eta$ learning rate $10^{-3},10^{-4},10^{-5}$] $10^{-4}$
$\omega$ weight decay $10^{-4},10^{-5},10^{-6}$] $10^{-5}$
@@ III.4 Training and testing
Figure 5: Loss function evolution per network over epochs. From top to bottom, represent the results of training for T, E, and B maps, respectively. The dark blue solid line indicates the training set loss function evolution and the light blue solid line indicates the validation set loss function evolution.
@@
UNet networks are predicated on fully convolutional networks, which comprise a convolutional network and an inverse convolutional network. Thus, the heart of these networks is the convolutional layer, which involves convolving filters on the input data.
@@
In alignment with the works of Makinen et al. and Ni et al. Makinen:2020gvh; Ni:2022kxn, we have fixed the number of convolutional kernels at the outset of the input to $32$. The kernel size determines the convolution’s field of view, which we have established at $3\times 3$. To achieve the requisite output dimensionality, we have employed “same” padding to manage sample boundaries in both convolutions and transpose convolutions. The stride parameter specifies the kernel’s traversal steps across the image. In our model, we have maintained the default settings of ${\rm stride}=1$ for convolutions and ${\rm stride}=2$ for transpose convolutions.
The UNet++ architecture was used to train the CMB delensing process via a set of lensed and unlensed CMB sky maps in an end-to-end fashion. Table 1 displays the hyperparameters’ specifics used in this network. The NAdam optimizer was utilized in the analysis with the standard TensorFlow parameters ruder2016overview; Sashank:2019abs.
@@
The hyperparameters were meticulously fine-tuned for network optimization. The batch size and the initial number of convolutional filters are optimized to $32$, $64$ and $128$, respectively, being restricted by GPU memory. The number of epochs is fixed to $3000$; Table 1 illustrates the learning rate setup. Meanwhile, weight decay was examined with a list of previous values listed in Table 1. The optimized values of the initial number of convolution filters, learning rate, weight decay, and batch size were fixed at $32$, $10^{-4}$, $10^{-5}$, $64$, respectively. The entire number of trainable parameters is $7.4\times 10^{7}$. The total number of trainable parameters is $9.04\times 10^{6}$. Since the CMB sky map data contains both positive and negative values due to temperature fluctuations, ReLU and LeakyReLU activation with alpha=1.0 was chosen to handle this relationship. Figure 5 illustrates the evolution of the loss function during the training of T, E, and B maps as a function of the number of epochs. The figure shows the training results of T, E, and B maps in a top-to-bottom order. The dark blue line indicates the evolution of the training set loss function, while the light blue dashed line represents the validation dataset’s loss function development.
In this study, we utilized a high-performance computing platform equipped with eight NVIDIA A40 GPUs, an Intel Xeon Platinum 8358 processor, and 1000 GB of memory, providing robust computational support for the training of complex models. Taking a set of parameter-generated sky maps (e.g., the T field) as an example, a single training run takes approximately 16 hours. For datasets containing all three fields (T, E, and B), after applying 30 different rotational transformations, a total of 90 training tasks were performed, accumulating to 1440 hours of total training time.
IV Results and Discussion
In this section, we conducted a thorough analysis of the results generated by the QE algorithm and the UNet++ model. Specifically, we systematically explored the outcomes of these two methods from two key perspectives: sky map patches and power spectra.
IV.1 Sky map analysis
Figure 6: Comparison of training and predicted sky map patches. The image is arranged in columns corresponding to the three sky map patches T, E, and B in order. The columns are arranged from top to bottom, with the first row shows the sky map patch that has been lensed and is affected by noise and instrumental effects. the second row showing the original sky patch of the real sky image that is not affected by gravitational lensing effects, the third row showing the lens-removed prediction result of the UNet++ model after processing, the fourth row showing the lens-removed prediction result of the QE algorithm after processing, the fifth row showing the residual map of the UNet++ prediction result and the real image, and the sixth row showing the residual map of the QE algorithm prediction result and the real image. Similar to Figure 3, the size of a sky patch is $214.86~{\rm deg}^{2}$, with the center located at $(l,b)=(78.75{\rm{}^{\circ}},0{\rm{}^{\circ}})$. The unit is $\rm\mu K$}.
@@ IV.1 Sky map analysis
Based on the network parameter configuration described in the previous chapter, we have trained multiple datasets and performed prediction operations on the test set. Figure 6 shows the residual performance between the predicted and true values of UNet++ model processing results. Firstly, the first and second rows of the modified figure display a sky map patch from the test set, which includes lensing effects, noise, and instrumental effects, alongside its corresponding label (an unlensed sky map patch). Secondly, the third and fourth rows present the delensing effect visualizations generated by the UNet++ and QE algorithms, respectively. Finally, the fifth and sixth rows illustrate the differences between the delensing results produced by the UNet++ and QE algorithms and their respective labels.
@@
SSIM is mainly used to analyze the structural information of two images wang2004image; wang2009mean. The closer the SSIM is to $1$, the more similar the structural information of the two images is. We have generally processed SSIM by setting the parameters $\alpha,\beta,\gamma$ to $1$. For the sky patches of two fields $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, we have carried out the following treatment:
$\begin{split}{\rm SSIM}(\mathcal{F}_{1},\mathcal{F}_{2})&=[l(\mathcal{F}_{1},\mathcal{F}_{2})^{\alpha}\cdot c(\mathcal{F}_{1},\mathcal{F}_{2})^{\beta}\cdot s(\mathcal{F}_{1},\mathcal{F}_{2})^{\gamma}]\\ &=\frac{(2\bar{\mathcal{F}}_{1}\bar{\mathcal{F}}_{2}+c_{1})[2{\rm CoV}({\mathcal{F}_{1},\mathcal{F}_{2}})+c_{2}]}{(\bar{\mathcal{F}}_{1}\bar{\mathcal{F}}_{2}^{2}+c_{1})[\sigma(\mathcal{F}_{1})^{2}+\sigma(\mathcal{F}_{1})^{2}+c_{2}]},\end{split}$ (17)
where $l(\mathcal{F}_{1},\mathcal{F}_{2})$, $c(\mathcal{F}_{1},\mathcal{F}_{2})$, $s(\mathcal{F}_{1},\mathcal{F}_{2})$ represent the brightness comparison, contrast comparison, and structure comparison between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, respectively. ${\rm CoV}(\mathcal{F}_{1},\mathcal{F}_{2})$ represents the covariance between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, while $\sigma(\mathcal{F}_{1})$ and $\sigma(\mathcal{F}_{2})$ represent the standard deviation of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, respectively. $c1$ and $c_{2}$ are constants to prevent the numerator or denominator from being zero.
According to Equation (17), we can give the value of SSIM, as shown in Table  2. From the table, it can be seen that the SSIM values of the T, E, and B sky map patches after delensing processing using the UNet++ algorithm are almost close to $1$. This means that the image blocks of the sky map after lens processing using the UNet++ algorithm are very similar in structure to those without lens.
In contrast, the results of the QE algorithm are much lower than those obtained by the UNet++ algorithm. Especially in the power spectrum of B-mode, this difference is more obvious. This can be explained by the fact that the UNet++ algorithm is better than the QE algorithm in removing lensing effects.
Table 2: SSIM values between predicted map patches and ground truth labels.
${\rm SSIM}_{({\rm T},{\rm T})}$ ${\rm SSIM}_{({\rm E},{\rm E})}$ ${\rm SSIM}_{({\rm B},{\rm B})}$
${\rm Patch}_{({\rm Truth},QE)}$ 0.7667 0.7390 0.0282
${\rm Patch}_{({\rm Truth},UNet++)}$ 0.9880 0.9760 0.9544
@@
$PSNR(\mathcal{F}_{1}-\mathcal{F}_{2})=10\cdot\log_{10}[\frac{{\rm MAX}(\mathcal{F}_{1},\mathcal{F}_{2})}{{\rm MSE}(\mathcal{F}_{1},\mathcal{F}_{2})}]{\rm dB},$ (18)
@@
where ${\rm MAX}(\mathcal{F}_{1},\mathcal{F}_{2})$ represents the maximum pixel difference between images $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, and ${\rm MSE}(\mathcal{F}_{1},\mathcal{F}_{2})$ represents the mean squared error between $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. Through this calculation, we can more comprehensively evaluate the degree of distortion at the pixel level in the image, and thus provide strong support for the comprehensive evaluation of image quality.
@@
PSNR stands as a simple yet widely utilized metric for assessing image quality, employing the decibel (dB) as its unit of measurement jahne2005digital; bovik2010handbook. Within this evaluative framework, a higher numerical value signifies a lower degree of image distortion, thereby indicating superior image quality. Akin to the computation of the SSIM metric, we have also conducted a corresponding analytical calculation of the PSNR, as shown in Table 3.
Table 3: PSNR values between predicted map patches and ground truth labels. The unit is $\rm dB$}.
${\rm PSNR}_{({\rm T},{\rm T})}$ ${\rm PSNR}_{({\rm E},{\rm E})}$ ${\rm PSNR}_{({\rm B},{\rm B})}$
${\rm Patch}_{({\rm Truth},QE)}$ 21.14 19.25 14.48
${\rm Patch}_{({\rm Truth},UNet++)}$ 37.70 38.76 37.87
By analyzing the data in Table 3, we can find that the PSNR ratio obtained by the UNet++ algorithm is also significantly higher than that of the QE algorithm. This result shows that the UNet++ algorithm performs better in signal recovery and can restore the true signal more effectively.
IV.2 Angular power spectrum analysis
In the previous data simulation stage, in order to study the impact of the lensing effect on the CMB, we conducted 30 simulations for each component. Considering that the segmentation operation of the sky map may introduce errors in scale information, in order to reduce the errors caused by this segmentation, we performed a specific angle rotation on the simulated sky map, followed by a delensing operation, and finally rotated back to the original angle. After such processing, we averaged the results to obtain a complete sky map for subsequent angular power spectrum calculations.
To more intuitively present the impact of rotation processing on the results, we first show the power spectra of each component throughout the day without rotation processing, as shown in Figure 7. In this figure, we present the TT, EE, and BB power spectra of the CMB without lensing effect processing. The temperature power spectrum is represented by a dark red solid line, which represents the angular power spectrum generated by the T signal without lens processing. The long red dashed line represents the angular power spectrum generated by the T signal without lens processing through the UNet++ network. The light red dotted line represents the angular power spectrum generated by the T signal without lens processing through the QE algorithm. The polarization power spectrum covers the EE and BB spectra, and is represented by dark blue, blue long dashed lines, light blue dotted lines, dark green, green long dashed lines, and light green dotted lines in a similar manner to represent the angular power spectra generated by the T signal without lens processing, UNet++ lens processing, and QE algorithm lens processing.
Figure 7: For the TT spectrum, the dark red solid line denotes the lensed power spectrum, the red dashed line shows the unlensed spectrum, the salmon dash-dot line corresponds to the spectrum delensed using UNet++, and the pink dotted line represents delensing via the QE algorithm. In the EE panel, the dark green solid line indicates the lensed $E$ power spectrum, the green dashed line shows the unlensed case, the light green dash-dot line represents delensing with UNet++, and the lime dotted line shows the delensed spectrum obtained by the QE method. For the BB spectrum, the dark blue solid line denotes the lensed BB power spectrum, the blue dashed line represents the unlensed primordial signal, the slight blue dash-dot line corresponds to delensing by UNet++, and the cyan dotted line shows the result from QE-based delensing. The parameters for the temperature power spectrum are set to $A_{s}=2.1\times 10^{-9}$ and $r=0.005$.
As can be seen from Figure 7, there are significant differences between the delensing effects of the UNet++ and QE algorithms at high $\ell$ values (small scales) and the label power spectrum. This phenomenon reveals the varying performance of the QE and UNet++ network across different scales. Specifically, as the scale decreases, the delensing effect of UNet++ diminishes gradually, and the rate of this decrease accelerates. We propose two possible explanations for this finding.
Firstly, we believe that the lensing effect may gradually increase on smaller scales, causing a significant distortion of the original CMB signal. This distortion may exceed the correction capability of the UNet++ network, leading to a decrease in delensing effectiveness on even smaller scales. Secondly, after dividing the full-sky map into small patches and performing delensing processing on them separately, the algorithms can reconstruct the central regions of each patch relatively accurately, but often fail to fully recover or show slight misalignments at the edges. After stitching the patches together, the tiny discontinuities at the boundaries of each patch correspond to the high $\ell$ components in the angular power spectrum. In contrast, low $\ell$ modes span multiple small patches, and the algorithms tend to retain the average values of each patch. Therefore, even if there are subtle offsets in the overall large - scale structure after stitching, it can still maintain good continuity. To reduce the impact of boundary artifacts on high $\ell$ signals, we attempt to perform rotation processing on the small patches.
It is worth noting that the two algorithms also exhibit opposite deviations in the high $\ell$ range. This is because UNet++, in order to suppress noise amplification, applies stronger smoothing or regularization to strong signals, resulting in a power spectrum recovered at high $\ell$ values that is lower than the label. On the other hand, the QE algorithm is more inclined to preserve all the details, leading to a power spectrum at high $\ell$ values that is higher than the label.
Figure 8: Angular power spectrum of the unlensed and predicted CMB $T$, EE and $B$. The results of UNet++ are obtained by averaging over 30 rotational transformations, while the power spectra of the QE algorithm are calculated based on the full-sky map. The line color and line style in this diagram are consistent with Figure 7.
To optimize the delensing results,as previously stated, we initially applied 30 rotational transformations to the acquired full-sky maps. Subsequently, we mitigated the lensing effects present in the rotated data. Finally, we applied another rotational operation to the processed data, culminating in the results depicted in Figure 8. Among them, the results obtained through the QE delensing algorithm are directly processed based on the full-sky map. Consequently, no additional rotation and segmentation operations are required for this result. As evident from the figure, the errors stemming from image segmentation have been substantially reduced following the rotational processing.